OpenStudy (anonymous):

28 men can finish the job in 60 days. At the start of the 16 day, 5 men were laid off and after the 45th, 10 more men were hired. How many days were delayed in finishing the job? Please explain, step by step.

OpenStudy (anonymous):

let eaceh man do x amount of work. so the total amount of work to be done is 28*x*60=1680x work done in first 15 days = 28*x*15 =420x work done in the next 30 days =23*x*30=690x remaining work = 1680-690-420=570x so remaining days = 570x/33x=17.27 so total days are 62.27 2.27 days excess

OpenStudy (anonymous):

a repeat yes?

OpenStudy (anonymous):

yes, i just wanted to see other possible solutions, but thanks to the both of you! :)